OpenStudy (idealist10):

Solve 6x^2*y^2*dx+4x^3*y*dy=0.

OpenStudy (idealist10):

@phi

OpenStudy (idealist10):

Do you know how to solve this exact equation?

OpenStudy (phi):

I would factor out 2 x^2 y from each term

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (phi):

are you saying you can't factor out 2 x^2 y ? to get 3 y dx + 2 x dy = 0 which is separable into \[ - \frac{3}{2} \frac{dx}{x} = \frac{dy}{y} \]

OpenStudy (idealist10):

But the answer in the book says 2x^3*y^2=C and I did it your way by making it separable and I didn't get the same answer in the book. Is the book's answer correct?

OpenStudy (phi):

take the square root of both sides \[ \sqrt{2} x^\frac{3}{2} y = \sqrt{C} \\ y = \frac{\sqrt{C}}{\sqrt{2} x^\frac{3}{2} } \\ y = C' x^{-\frac{3}{2} } \]

OpenStudy (phi):

\[ - \frac{3}{2}\int \frac{dx}{x} = \int \frac{dy}{y} \\ - \frac{3}{2}\ln(x) + C = \ln(y) \\ \ln(y) = \ln(x^{- \frac{3}{2}}+C ) \] make each side the exponent of e \[ y = e^C x^{- \frac{3}{2}} \\ y = C' x^{- \frac{3}{2}} \]

OpenStudy (phi):

** paren put in the wrong spot for \[ \ln(y) = \ln(x^{-\frac{3}{2}}) + C \]

OpenStudy (phi):

Here is how they might want you to solve this. if you have P dx + Q dy First, show that you have an exact differential d/dy P = d/dx Q P = 6x^2 y^2 d/dy P = 12 x^2 y Q= 4 x^3 y d/dx Q = 12 x^2 y they are equal so it's exact. we now integrate using the formula \[ f(x,y) = \int_{x_0}^x P(x,y) dx + \int_{y_0}^y Q(x_0,y) \ dy = c \] where \( (x_0, y_0) \) is a point in the region where the integral exists. In this case (0,0) works (and is the simplest point to use) in particular, Q(0,y) is 12 0^2 y = 0. i.e. we only have one integral to do \[ f(x,y) = \int_{0}^x 6x^2 y^2 dx = C\\ 2 x^3 y^2 = C \] which is probably how the text wants you to solve the problem.

OpenStudy (idealist10):

\[2x ^{3}y ^{2}+g(y)\]